OpenStudy (anonymous):

What is the first step in solving this system of equations by elimination? 5a + 3b = –9 2a – 5b = –16 A. Subtract the two equations. B. Add the two equations. C. Solve for a. D. Multiply the first equation by 2 and the second by –5.

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 5a + 3b = -9&{\color{red}{ \times 5}}\implies &25a\cancel{+15b}=-45\\ 2a - 5b = -16&{\color{red}{ \times 3}}\implies &6a\cancel{-15b}=-48\\ \hline\\ &&31a\qquad \qquad=-93 \end{array}\) then solve for "a" once you find "a", substitute that on either equation and solve for "b" notice the step, so how we "eliminated" the " b's "